OpenStudy (anonymous):

Find the magnitude and direction angle of the vector v. v=2i + 5j

OpenStudy (anonymous):

euclidean geometry?

jimthompson5910 (jim_thompson5910):

|| v || = sqrt(a^2 + b^2) || v || = sqrt(2^2 + 5^2) || v || = sqrt(4 + 25) || v || = sqrt(29) --------------------------------- theta = arctan(b/a) theta = arctan(5/2) theta = 68.19859

OpenStudy (anonymous):

thanks!!!!!! i did get it right!! :)